A basketball is shot from an initial height of 2.40m with an initial speed of 15m/s directed at an angle of 35 degrees above the horizontal. How far from the basket (in meters) was the player if he made a basket
1
Expert's answer
2021-10-12T16:23:18-0400

The range is


"R=vT\cos\theta."

The time of flight upward:


"t_1=\frac {v\sin\theta}g."

The time downward:


"t_2=\sqrt{\frac{2h}g}=\sqrt{\frac{2(v^2\sin^2\theta/(2g)+h_0)}g}."

Range:


"R=v\cos\theta(t_1+t_2)=\\\space\\
=v\cos\theta\bigg(\frac {v\sin\theta}g+\sqrt{\frac{2(v^2\sin^2\theta/(2g)+h_0)}g}\bigg)=\\\space\\
=24.58\text{ m}."